OpenStudy (anonymous):

Identify the 31st term of an arithmetic sequence where a1 = 26 and a22 = -226.

sam (.sam.):

a22 = -226 a1+(n-1)d=-226 26+(22-1)d=-226 d=-12 \[T_{31}=a+(n-1)d\] 26+(31-1)(-12)=-334

OpenStudy (anonymous):

wb this one identify the 27th term of an arithmetic sequence where a1 = 38 and a17 = -74.

OpenStudy (anonymous):

Identify the 42nd term of an arithmetic sequence where a1 = -12 and a27 = 66